It takes a passenger train 3 hr less time than it takes a freight train to make the trip from Hugesville to boggstown. The passenger train averages 64km/h, while the freight train averages 32km/h. Find the distance from Hughesville to Boggstown. How far is it from Hughesville to Boggstown?
Answer by lwsshak3(11628)

It takes a passenger train 3 hr less time than it takes a freight train to make the trip from Hugesville to boggstown. The passenger train averages 64km/h, while the freight train averages 32km/h. Find the distance from Hughesville to Boggstown. How far is it from Hughesville to Boggstown?
***
let x=travel time of freight train
x-3=travel time of passenger train
speed*travel time=distance
..
64(x-3)=32x
64x-192=32x
32x=192
x=6
32x=192
How far is it from Hughesville to Boggstown? 192 miles